OpenStudy (lena772):

A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x).

OpenStudy (lena772):

Function of Stock D is f(x) = 23(1.07)^x

OpenStudy (lena772):

Gordon is evaluating stocks...

OpenStudy (the_fizicx99):

You're adding g(x) to your f(x) function so it's f(x) + g(x) = 23(1.07)^x - 6

OpenStudy (lena772):

Thanks

OpenStudy (the_fizicx99):

Yw